50 recent changes in FLAO Web retrieved at 21:33 (GMT)

FLAO2TelLog20160209
Allineamento HP con retro SX: x y z Rx Ry Rz 1.37 0.14 3.95 240 208 0 NCPA Disturbi: Seeing TN 0.0 165645 0.6 1650...
ConFiguraTions
Configurations FLAO#2 FLAO#2 @ Left Rear Bent Gregorian M3 position Official position, confirmed 2012 02 Tip Tilt Piston Sel Date Note 1115 ...
FLAO2TelLog20160208
Today (Feb 8th) AGW is reinstalled after the W board shimming done in clean room. Some problems with the telescope delay things, and we have the complete system (...
LbtReview
LBT AO review Date: 30th and 31st March 2009 Agenda: * FLAO_review_AgendaV3.pdf: FLAO review agenda Presentations: * FLAO_01.ppt: F. Quiros: FLAO system ...
r79 - 09 Feb 2016 - 01:28 by AoAdmin
LabLog20160207
IRTC removal After shimming is completed the IRTC is decoupled from the AGW unit and stored away Alignment verification We double check the small bar mirror, to ...
WActivityInTelescopeCleanRoom
* LabLog20100215 * LabLog20100216 * LabLog20100217 * LabLog20100218 * LabLog20100219 * LabLog20100220 * LabLog20102222 * LabLog20102223 Snow ...
LabLog20160206
AOS problem While setting up for adsec DX measurement, we find that the AOS Right reports "FLAO disconnected" and does not send elevation data, even if from the...
TelLogLBT672a20160206a
Main.RunaBriguglio 06 Feb 2016 Alignment alignment lost: new alignment: 1.1 1.92 3.4 87 136 fsm_load_shape(myflat, force_max_mode=200); myflat=20160205_00542...
LabLog20160205
New functional test document Following the functional checks done in the previous days, a new version (revision "C") of the W unit functional check document is up...
TelLogLBT672a20160205a
Main.RunaBriguglio 05 Feb 2016 Measurement of ff matrix first test, 2 iterations 20160205_140944 data seem good, no bad actuators final test 10 iterations 2016...
TelescopeLogLBT672a
LBT672a related Logs Mar 2010 Log. First Seeing Limited Run * TelLog672a20100311 * TelLogLBT672a20100312 First night * TelLogLBT672a20100312Day * Tel...
LabLog20160204
Cooling situation WFS starts up fine. Cooling line is closed because yesterday's leak has not been fixed, so we can keep the unit on for just short intervals. Pu...
TelLogLBT672a20160204a
Main.RunaBriguglio 04 Feb 2016 collect_ff_data da riga di comando (linea per linea) startup problems at the initial alignment new alignment: 1.94 0.02 3 218 2...
TelLogLBT672a20160203a
Preliminary Alignment M3: 1150 750 0 Sel = 38.85 hexapod: 2 0 1 150 450 0 Shell set successfully the 4D image is highly saturated, although the gain i...
LabLog20160203
Clean room Greg, Alfio We find W#1 in clean room ready for testing. We start attaching the network (fiber Mini Mc spare ethernet cable port 19 on switc...
TelLogLBT672a20160201a
Main.RunaBriguglio 01 Feb 2016 Preliminary alignment with new 4D focal station M3:
SHARKNightTest20150604
Fernando (Mountain), Mauro, Emanuele G. (Roma), Alfio (Arcetri) Aspettiamo fino alle 2 AM perche' prima c''e LBTI e poi il seeing e' cattivo (1.5") Albireo, mag=...
r3 - 13 Jan 2016 - 09:42 by EnricoPinna
PortingToQt4
Porting to Qt4 Some progress has been made in porting GUIs to Qt4. However, the task looks too big to be completed in time for the UAO release. Changeset with com...
RemoteAccess
Remote access to AO servers @ LBTO Simple access To access wfsdx server remotely (from outside LBTO): * Connect to ssh bridge: ssh X C c blowfish X userid...
FLAO1LabLog20130127
Ghiaccio e neve, lavoriamo con il retro Juan Carlos e Julin riallineano il sistema (usando la procedura)con il terziario nela posizione trovata di notte da J. Hil...
SHARKSimulData20151028
Simulation data for Padova group: phase residuals at different magnitudes and seeing.This time we have very long data cubes (100 seconds of simulated time) to ena...
SHARKSimulData20150701
Simulation data for Padova group: phase residuals at different magnitudes and seeing. Tracknum Mag Seeing SR Time (s) 20150701_143246 10 0.80 ...
FLAO1TelLog20141102
h7:30 Retro ON Abbiamo homato e riposizionato esapodo e terziario ma lo spot non torna nel W. AO print, hexa_spiral(30,15,/apply) Andiamo a controlare le viti de...
r9 - 29 Sep 2015 - 14:03 by EnricoPinna
LabLog20140731b
Setup In Zemax we represented a symplified version of the FLAO and LUCI optical systems: 1 The Pyramid arm of the FLAO WFS up to the detector pupil plane (w/o ...
CentOS6xFLAO
Test of FLAO Supervisor on CentOS 6.x servers. Resume The test has been performed by A.Puglisi and L.Fini on dec. 27/28 2012 using two spare servers at LBTO with...
r4 - 25 Sep 2015 - 07:51 by AoAdmin
Fsm20CmdReview
AdSec commands Cmd name Used by: ArbIntf ICE intf FSM2.0 name Notes RequestAdSecStatus Gui X GetAdsecStatus LoadProgram Gui, AOArb...
ScheDule
Main.EnricoPinna 22 Sep 2015 * SM20_schedule_v01.png:
NEW - 22 Sep 2015 - 08:10 by EnricoPinna
FlaoSimulator
FLAO WFS simulator Setup from SVN * FlaoSimulatorFromSVN or Download: * flao.tar.gz WFS simulator * flao_adsec.tar.gz Adsec simulator * hosts...
HowToPeerMsgDs
How to setup MsgD peering Message daemons can communicate between them using a custom "peering" mechanism: each MsgD is identified by a string (called "identity")...
TcsTelemetry
Enabling and usingTCS Telemetry NOTE: TCS dismissed mysql based telemetry implementation and is now writing to HDF files onto /lbt/telemetry_data. The following i...
FlaoSimulatorInstructions
1. Extract the tarball: $ tar xzvf flao.tar.gz $ cd flao 2. Update the computer's /etc/hosts file with the list of hostnames found in the "hosts" file (the list...
Fsm20TaskList
Upgrade of existing FLAO system with a new, more flexible AO Arbitrator and a new operator GUI. Schedule: Project file PNG file * New sequence definition ...
Fsm20IceInterface
Two ICE interfaces for the new AO Arbitrator: * aoArb.ice: interface exposed to the AOS * aoArbEng.ice: interface for the new operator GUI. Main.AlfioPugl...
Fsm20NewSequence
Word file with new sequence description and GUI mock up: * AO_supervisor_FSM2.0_draft_V04.doc: Main.AlfioPuglisi 31 Aug 2015
LabLog20150823
Sostituzione cube stage Laser di Argos in fiorriera e spariamo il laser riflesso sul soffito. In posizione W_internal il cubo crea uno spot sul soffitto a 290cm d...
r3 - 23 Aug 2015 - 23:45 by EnricoPinna
LabLog20150822
Diagnostic cube rotator stage Proviamo a rimisurare la strana impedenza fra Motor vs GND e Motor VS gnd. Al DB15 dello stage installato, se alimentato con i ...
NEW - 23 Aug 2015 - 03:42 by EnricoPinna
AgwControlSoftware
Main.AndreaTozzi 29 Aug 2007
LabLog20150821
J. Storm, E. Pinnna, M. Lefebrve Cable routing of the TT ARGOS unit Passiamo il bundle di fibre della QC e il cavo dell'aperture wheel. Facciamo un check ksu tut...
r3 - 22 Aug 2015 - 00:13 by EnricoPinna
LabLog20150820
Check visuale posizione del laser sull'aperture wheel FW#1 R90T10FW#2 controled by the aperture wheel controller ADC1 0ADC2 0X= 74.95mm; Y= 51.40mm; Z=3...
r3 - 21 Aug 2015 - 01:16 by EnricoPinna
LabLog20150819
Ieri usando il sistema ogni poco ci si doveva fermare a causa della temp della stratiz della BCU39 che sia vvicinava a 60C. Raffreddamento lento e perdita di temp...
r5 - 20 Aug 2015 - 02:22 by EnricoPinna
LabLog20150817
M. Bonaglia, E. Pinna, J. Storm, G. Rahmer 08:45 W#1 Functional check Stephen Hooper ha finalmente abilitato la porta 21 sullo switch104 per il W#1 (subnet 192.1...
r6 - 19 Aug 2015 - 15:37 by EnricoPinna
LabLog20150818
Check filter wheel Ripartiamo col sistema e ora l'homing della FW#2 รจ di nuovo con la QC in pos=0. Allineamento QC La migliore posizione per centraggio dello spo...
r4 - 19 Aug 2015 - 01:25 by EnricoPinna
LabLog20140731
h08:30 Check FW#2 su W#1 La FW fa ancora dei rari episodi di stop toccando il connettore sulla FW e i suoi fili adiacenti. Sembra ceh siano i fili che vanno al co...
LabLog20150816
M. Bonaglia, E. Pinna, J. Storm, G. Rahmer 16:30 W#1 connection.... W#1 in electronic lab. Collegiamo la rete al solito cavo ethernet collegato al SW103 port 15....
r2 - 17 Aug 2015 - 01:18 by EnricoPinna
LbtTel
Logs and notes of FLAO #1 and #2 Commissioning after mounting on telescope Notes about the system after October 2011 run Oct2011Notes Software TODO list @ teles...
SHARKDayTest20140604
NCPA measuerement senza avere delle slopenull inserite per sbaglio Ripartiamo dal set di NCPA ottenute ieri. Scan automatico con SR misurato su psf Shift Add Z4fu...
SHARKNightDay20150603
Trasferimento automatico immagini da SHARK a elablib. PSF VS FoV position Centrale TN150719 TN151237 X= 0.15mm TN151319 X= 0.15mm TN151432 Y= 0.15mm TN151513 Y...
NEW - 03 Jun 2015 - 23:49 by EnricoPinna
SHARKDayTest20150601
SHARK day test: Fernando P. (Mountain), Mauro (remoto da Roma), Alfio (remoto da Arcetri) h7:00 (Tucson time) lavorano sul PC di SHARK Stanno spostando il PC di ...
SHARKNightTest20150602
Night test: Fernando P. (Mountain), Vanessa (Tucson remote) Mauro (Roma), Alfio (Arcetri) Simone h00:30 (Tucson time) telescopio HIP 85313 mag=7.3 Seeing 1.9", ...
FLAO1TelLog20150530
30 maggio: Doug (remote control room) Alfio (Arcetri) Jochen, Dave, Steve et al. (Mountain) Sommario Circa 3:30 ore in cielo. Condizioni generalmente buone: seei...
Number of topics: 50
< Previous Page 2 of 28 Next >

See also: rss-small RSS feed, recent changes with 50, 100, 200, 500, 1000 topics, all changes
Topic revision: r1 - 26 Feb 2018, ProjectContributor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback